A 0.750-cm diameter shell contains a uniformly distributed charge of 65.4 pC on its surface. Determine

the potential near the surface of the sphere.


1
Expert's answer
2022-02-05T10:28:09-0500

Given:

"d=0.750\;\rm cm"

"q=65.4*10^{-12}\;\rm C"


The potential near the surface of the sphere is given by

"V=k\frac{q}{R}"

"V=9*10^9\frac{65.4*10^{-12}}{0.750/2*10^{-2}}=157\:\rm V"